The sum of Megan, Nicholas and John's age is 51. Megan is 5 years older than Nicholas and Nicholas is 8 years older than John. What is each of their ages?
Answer by Alan3354(69443)

The sum of Megan, Nicholas and John's age is 51. Megan is 5 years older than Nicholas and Nicholas is 8 years older than John. What is each of their ages?
-----------------
M + N + J = 51
M = N + 5
N = J + 8
etc